OpenStudy (anonymous):

How do you find the quadratic equation using matrices with the given points (3,91) (-1,27) (4,112)

OpenStudy (anonymous):

@Hero i know but the teacher wants us to use matrices

hero (hero):

The general form of a quadratic equation is y = ax^2 + bx + c

hero (hero):

Inputting each point in to the equation gives the following system: 91 = 9a + 3b + c 27 = a - b + c 112 = 16a + 4b + c

OpenStudy (anonymous):

what do you do after that then?

hero (hero):

First you'll have to rewrite the system like so 9a + 3b + c = 91 a - b + c = 27 16a + 4b + c = 112 Then treat a, b, c as the variables and the numbers as the coefficients.

OpenStudy (anonymous):

this whole concept is very confusing to me. This extra credit assignment is worth about half of a test so i really want to get this right.

hero (hero):

Then you can setup the matrix. Only thing is, you'll run into a problem while solving.

OpenStudy (anonymous):

what problem?

hero (hero):

Actually, they probably mean for you to solve this using Cramer's Rule like that link @zzr0ck3r provided

OpenStudy (anonymous):

oh okay. Thanks then!

hero (hero):

Someone posted this same exact question about a week ago and no one could figure it out. They said "Solve using matrices" instead of "Solve using Cramer's Rule" In a way, it's a trick question, but not really.

OpenStudy (anonymous):

Just a different name? Probably the same class as me haha.

hero (hero):

When a question says "Solve using matrices", Cramer's Rule isn't the first thing that comes to mind.

hero (hero):

But from now on, you can be on the lookout for such questions.

OpenStudy (anonymous):

Oh. What?

hero (hero):

A large percentage of math questions are structured this tricky way where it seems unfamiliar, but with a bit more creative thought, you can think of the right approach to answering it.

OpenStudy (anonymous):

THANK YOU MAN

hero (hero):

@jayz657, do you have anything you wish to contribute?

OpenStudy (anonymous):

oh sorry i was afk from open study but nope everything is all good :)